OpenStudy (ttop0816):

Is this right??? Position vectors m and n have terminal points of (2, 3/4) and (1, y), respectively. If m · n = 8, then y is _____. 4 1/2 6 1/2 <<<<<<<<<<<< is this correct?? 8 12

OpenStudy (ttop0816):

@agent0smith please check?

OpenStudy (ttop0816):

oh this is 8

OpenStudy (agent0smith):

Use the dot product: m*n = 2*1 + (3/4)*y = 8

OpenStudy (agent0smith):

solve for y